OpenStudy (anonymous):

Find a positive number such that the sum of r times its reciprocal and t times the number is as small as possible.

OpenStudy (ilovecake):

1/x

OpenStudy (anonymous):

im using online homework and i put that in and it didnt work says wrong still :/

OpenStudy (ilovecake):

:(

OpenStudy (ranga):

Let the number be x and let the sum be y. Sum y = "r times its reciprocal and t times the number" y = r/x + tx Sum should be as small as possible. To minimize y, find dy/dx, equate to zero and solve for x. (r and t are constants).

OpenStudy (anonymous):

still cant get it lol to many variables

OpenStudy (ranga):

You can treat r and t as constants. y = r/x + tx = rx^-1 + tx Find the derivative dy/dx = -rx^-2 + t Equate dy/dx to zero and solve for x. -r/x^2 + t = 0 t = r/x^2 x^2 = r/t x = sqrt(r/t)

OpenStudy (anonymous):

thanks man :)

OpenStudy (ranga):

\[\Large x = \sqrt{\frac{ r }{ t }}\]

OpenStudy (ranga):

You are welcome.